PLEASE HELP WITH ENTIRE EXERCISE

A) You invested $50,000 in a mutual fund at the beginning of the year when the NAV was $68.23. At the end of the year, the fund paid $0.32 in short-term distributions and $0.49 in long-term distributions. If the NAV of the fund at the end of the year was $72.59, what was your return for the year?

B) A closed-end fund has total assets of $510 million and liabilities of $290,000. Currently, 29 million shares are outstanding.A mutual fund sold $148 million of assets during the year and purchased $118 million in assets. If the average daily assets oAn open-end mutual fund has the following stocks: Stock Α. Shares 15,500 31,000 7,500 73,000 Stock Price $ 79 29 57 D 15 The
Answer #1

A)the formula Total reruen se calculated by (Dt (NAV and NAVstay) x 100 NAV Start Hele D = Divideud declared for the - NAV StarB)NAV = HR Assers - liabilities Мед а акацdRusa 510,000 000 - о, ооо 24, Ооо, ооо = _ ~ $ 1. 5+62. 54 $ (т. 58 Price = Share NJAssets Assets Average sold = $ 148 million rucuased $118 million daily assets- $456 million Funds twinouee Mano - of MeulinumNAV= Assers - liabilities Number of outstanding herits Asseos - 15,500 x 79 + 31,000 + 2a + 7500 x5t 73,000 X 15 t = 1,22u, o